Three boys and three girls are to sit on a bench for a photograph. Find the number of ways this can be done if three girls must all sit apart.
Answer by Edwin McCravy(20064)

There are only these 4 ways to arrange them sex-wise:

GBGBGB, BGBGBG, GBBGBG, GBGBBG

For each of these 4 sex-wise arrangements, there are

3! ways to arrange the 3 girls, and
3! ways to arrange the 3 boys.

That's 4*3!*3! = 4*6*6 = 144 ways.

Edwin